# A Comparative Study on Enhancing Prediction in Social Network Advertisement through Data Augmentation

### v3

@@ I Introduction
The underlying hypothesis of this study is that synthetic data augmentation can serve as an effective solution to overcome the limitations of datasets. By generating new data points that mimic real data, the model can potentially learn more generalizable patterns, thereby improving accuracy when making predictions on unseen data. Similar methods have been taken in other fields such as CT segmentations  [5, 9, 10, 11], natural language processing [12], real-time systems [13, 14, 15], etc. Hence, our study endeavors to translate successful methodologies from other domains to the realm of social network advertising.
@@
How does synthetic data using GAN, VAE, and GMM affect the performance of various classifiers in the field of social network advertising?
Out of these data augmentation techniques, which one provides the best improvement in forecast accuracy?
@@ II-D Classification Model Training
Dense neural network: We build a multi-layer linear network, use the ReLU activation function, and train using the Adam optimization method. Similar settings have been used in previous research [16, 20, 21, 22].
@@
Logistic regression: We adjust regularization strength and solver type to handle nonlinear relationships in our data. This model is widely used in various fields [23, 24].
@@ V-B Theoretical and Practical Implications
The theoretical implications of these findings provide empirical support for using VAEs and GANs to create synthetic datasets that help build more accurate predictive models [29, 30, 31, 32, 33, 34, 35, 36]. In practice, these findings could influence how data scientists in the field of social network advertising build datasets, balancing the trade-offs between data privacy and the need for large-scale, balanced data.

### v4

@@ I Introduction
The underlying hypothesis of this study is that synthetic data augmentation can serve as an effective solution to overcome the limitations of datasets. By generating new data points that mimic real data, the model can potentially learn more generalizable patterns, thereby improving accuracy when making predictions on unseen data. Similar methods have been taken in other fields such as CT segmentations  [5, 9], natural language processing [10], real-time systems [11, 12], etc. Hence, our study endeavors to translate successful methodologies from other domains to the realm of social network advertising.
@@
How does synthetic data using GAN, VAE, and GMM affect the performance of various classifiers in the field of social network advertising.
Out of these data augmentation techniques, which one provides the best improvement in forecast accuracy.
@@ II-D Classification Model Training
Dense neural network: We build a multi-layer linear network, use the ReLU activation function, and train using the Adam optimization method. Similar settings have been used in previous research [13, 17, 18].
@@
Logistic regression: We adjust regularization strength and solver type to handle nonlinear relationships in our data. This model is widely used in various fields.
@@ V-B Theoretical and Practical Implications
The theoretical implications of these findings provide empirical support for using VAEs and GANs to create synthetic datasets that help build more accurate predictive models [23, 24, 25, 26]. In practice, these findings could influence how data scientists in the field of social network advertising build datasets, balancing the trade-offs between data privacy and the need for large-scale, balanced data.